# Gradient is All You Need: Gradient-Based Attention Fusion for Infrared Small Target Detection

### v1

@@ Abstract
Infrared small target detection (IRSTD) is widely used in civilian and military applications. However, IRSTD encounters several challenges, including the tendency for small and dim targets to be obscured by complex backgrounds. To address this issue, we propose the Gradient Network (GaNet), which aims to extract and preserve edge and gradient information of small targets. GaNet employs the Gradient Transformer (GradFormer) module, simulating central difference convolutions (CDC) to extract and integrate gradient features with deeper features. Furthermore, we propose a global feature extraction model (GFEM) that offers a comprehensive perspective to prevent the network from focusing solely on details while neglecting the background information. We compare the network with state-of-the-art (SOTA) approaches, and the results demonstrate that our method performs effectively. Our source code is available at https://github.com/greekinRoma/Gradient-Transformer.
@@
Infrared small target detection (IRSTD), convolution neural network (CNN), gradient Transformer, global feature extraction.
@@ I Introduction
INFRARED small target detection (IRSTD) is critically important for numerous space-based computer vision applications, including remote sensing [1], precision guidance [2], and marine rescue [3]. While IRSTD has been extensively studied, it remains challenging. The considerable imaging distances associated with space-based infrared cameras often result in small targets that lack detailed texture and shape information.
@@
Currently, the IRSTD methods can be categorized into model-driven and data-driven methods. Model-driven methods include three main approaches. 1) Filter-based methods, such as the Tophat [4] and Max-Median filters [5], enhance the target visibility in complex backgrounds while suppressing background noise. 2) Human visual system (HVS)-based methods, such as the local contrast measure (LCM)[6] and multi-patch contrast measure (MPCM)[7], leverage human visual perception to enhance local information. 3) Low-rank matrix decomposition and reconstruction methods, such as infrared patch-image (IPI)[8] and reweight image patch tensor (RIPT)[9], decompose small targets and backgrounds using low-rank matrices. However, these techniques require manual parameter settings and often need help to adapt to complex scenarios, particularly in high noise levels or significant background clutter.
@@
With the advancement of deep learning, data-driven approaches have achieved substantial progress in infrared small target detection (IRSTD). For example, the Asymmetric Context Modulation (ACM) network [10] introduces an asymmetric feature fusion technique, an alternative to the conventional skip connections in U-Net. Building on this, DNANet [11] implements a multi-layer nested architecture that supports progressive and adaptive interactions among feature layers. Moreover, UIUNet [12] enhances the detection of local target contrasts by integrating multiple U-Net structures and utilizing interactive cross-attention mechanisms for feature fusion. Additionally, GSTUnet [13] merges vision Transformer technology with CNNs in the encoder to efficiently extract edge features from small targets. Generally, these data-driven methods surpass traditional model-driven approaches. However, they have drawbacks, such as the substantial number of parameters and the potential loss of local feature information in higher layers, indicating opportunities for further enhancements.
In order to address the challenges of feature extraction, we propose the Gradient Attention Network (GaNet). GaNet incorporates two key innovation components to improve feature extraction. First, we propose a gradient Transformer (GradFormer) module that performs the feature gradient calculations by simulating the central difference convolutions (CDC) and enhances weak features to enable the network to learn a comprehensive feature representation of the target. Second, the global feature extraction module (GFEM) solves the problem of lacking global background perception, thereby improving the ability to obtain contextual information. The main contributions of this letter are as follows.
@@
    We propose the GANet, which utilizes the GradFormer to simulate the CDC with dynamic weights, extracting the edge and point information vital for IRSTD.
@@
    We conduct comparative experiments on the IRSTD-1K and NUDT-SIRST datasets, and the results show that GaNet outperforms existing methods.
@@ II Methods
The overall architecture of GaNet is illustrated in Fig. 1. GradFormer is utilized to extract gradient information, and GFEM enhances contextual features. Finally, detailed information from low-level layers is integrated with high-level semantics with the global receptive field.
Figure 1: The overall structure of GaNet. The blue represents the downsampling module, the green indicates the GradFormer module, the orange corresponds to the upsampling module, the red signifies the concatenate operation, and the purple denotes the global feature extraction module (GFEM).
II-A Gradient Transformer (GradFormer)
CDCs are defined in Eq. 1 and are employed to extract edge and gradient features. Here, $c_{i}$ represents the number of input channels, $o_{ixy}$ denotes the input value at the $i$-th channel and position $(x,y)$, and $b_{ijxy}$ refers to the $j$-th surrounding value around $o_{ixy}$. While CDCs excel at extracting details, there is still room for improvement. A significant limitation is the use of fixed weights for different directions, meaning the importance of each direction remains constant across different images.
$out_{\hat{c}xy}=\sum^{c_{i}-1}_{i=0}\sum^{7}_{j=0}w_{\hat{c}ijxy}(b_{ijxy}-o_{ixy})\$ (1)
We can transform Eq. 1 into a fusion of Eq. 2 and Eq. 3. The process described in Eq. 2 is illustrated in Fig. 2. As shown in Fig. 2, we use edge convolution to extract the difference values from the input $T\in\mathbb{R}^{c_{i}\times w\times h}$ and reshape them to produce a local difference matrix $D\in\mathbb{R}^{8c_{i}\times wh}$. In Eq. 3, we utilize the weight matrix $M_{w}\in\mathbb{R}^{c_{o}\times 8c_{i}}$ and the local difference matrix $D$ to obtain the final result $Out$. Here, $w$ denotes the width of the input, $h$ represents the input height, and $c_{o}$ signifies the number of output channels.
@@ II-A Gradient Transformer (GradFormer)
Figure 2: $D$ is obtained by performing edge detection convolution on $T_{i}$ and unfolding the output.
When $M_{w}$ is influenced by the input $T_{i}$, the CDC could exhibit enhanced robustness. Based on the analysis above, we introduce the GradFormer to simulate CDC and extract the details using dynamic weights from low-level layers for IRSTD. The structure of GradFormer is shown in Fig. 3.
Figure 3: The various heads in GradFormer employ the edge convolution to get $D_{\sum}$. Each head’s key and value are derived from $D_{\sum}$, while the query is generated from the flattened input.
In varied heads of the GradFormer, we employ different unfolding methods based on edge detection convolutions in Fig. 2 with varying dilation and reshape operations to get the unfolding .
In the various heads of GradFormer, we apply different edge detection convolutions from Fig. 2 to $T$, utilizing varying dilation ratios. The outputs from these convolutions are then reshaped into $D^{n}\in\mathbb{R}^{8c_{i}\times wh}$. Additionally, the input $T$ is reshaped to form $O\in\mathbb{R}^{c_{i}\times wh}$. The token $O$ serves as the query, and the local difference matrix $D^{n}\in\mathbb{R}^{8c_{i}\times wh}$ functions as both the key and the value.
@@
where $W^{n}_{Q}\in\mathbb{R}^{\frac{c_{o}}{m}\times c_{i}}$, $W^{n}_{K}\in\mathbb{R}^{\frac{8c_{i}}{m}\times 8c_{i}}$, $W^{n}_{V}\in\mathbb{R}^{\frac{8c_{i}}{m}\times 8c_{i}}$ are the reshaped weights of convolutions, $n$ denotes dilation ratios, and $m$ denotes the number of heads. We utilize the $Q^{n}\in\mathbb{R}^{\frac{c_{o}}{m}\times wh}$ and $K^{n}\in\mathbb{R}^{\frac{8c_{i}}{m}\times wh}$ to produce the attention matrix $M^{n}\in\mathbb{R}^{\frac{c_{o}}{m}\times\frac{8c_{i}}{m}}$. $V^{n}\in\mathbb{R}^{\frac{8c_{i}}{m}\times wh}$ denotes the value of the transformer. The result $Out^{n}\in\mathbb{R}^{\frac{c_{o}}{m}\times wh}$ is defined as follows:
@@
where $M^{n}_{mix}\in\mathbb{R}^{\frac{c_{o}}{m}\times 8c_{i}}$ denotes the dynamic matrix derived from a attention matrix $M^{n}$ and a fixed convolution weight matrix $W_{V}^{n}$. According to Eq. 3, $Out^{n}$ is the equivalent result of the CDC with dynamic weights and a dilation ratio of n. $Norm$ is the Instance Normalization. At last, we concatenate the results in varied heads, reshape them and use $1\times 1$ convolution to fuse them as shown in the Equation below:
@@
II-B Global Feature Extraction Module
Background information is crucial alongside detailed features in IRSTD. However, perceiving background relies on global features, which the GradFormer struggles to handle effectively. To address this limitation and enhance global information perception, we propose the Global Feature Enhancement Module (GFEM), as illustrated in Fig. 4.
@@ II-B Global Feature Extraction Module
Figure 4: Diagram of the Global Feature Extraction Module (GFEM) using the Non-local Attention Module to capture global spatial features and Squeeze-and-Excitation Block to extract global channel information.
@@ II-C Loss Function
The loss function utilized in our model training is the soft intersection over union (SoftIoU) loss, where the predictions of GaNet are mapped into the range of 0 and 1, using the Sigmoid function as shown in Eq. 7.
@@
$SoftIoU=1-\frac{\sum_{i,j}p_{i,j}\cdot g_{i,j}}{\sum_{i,j}p_{i,j}+\sum_{i,j}g_{i,j}-\sum_{i,j}p_{i,j}\cdot g_{i,j}}$ (8)
$SoftIoU_{loss}=1-SoftIoU$ (9)
@@
Figure 5: Visual results from varied data-driven methods. Red, green, and yellow boxes label detected targets, missed targets, and false alarms.
@@ III Experiment and Analysis
This section describes the experimental details, including a series of ablation studies to verify the proposed modules. Comparative experiments with other methods demonstrate that our approach outperforms existing state-of-the-art (SOTA) methods.
@@ III-A Dataset and Evaluation Metrics
The datasets used for evaluating the GaNet are NUDT-SIRST[11] and IRSTD-1k[16], which are widely employed to assess model performance. The evaluation metrics include mean intersection over union ($mIoU$), F-measure ($F_{1}$), the probability of detection ($P_{d}$), and false alarm rate ($F_{a}$).
@@ III-B Experimental Environment and Parameter Settings
Our model is implemented by the PyTorch framework on NVIDIA GeForce RTX 4080 GPUs. We employ the Adam optimizer, starting with an initial learning rate of $5\times 10^{-4}$. This learning rate is reduced to $5\times 10^{-5}$ at epoch 200 and further decrease to $5\times 10^{-6}$ We use a batch size of 4 for training and run the model for 400 epochs.
@@ III-C Ablation Study
TABLE I: Studies on different dilation rates on NUDT-SIRST (The best results are bold)
@@
TABLE II: Studies Of Different Components on NUDT-SIRST (The best results are bold)
Module mIoU(%) $F_{1}(\%)$ $P_{d}(\%)$ $F_{a}(10^{-6})$
U-Net 91.31 95.44 97.98 4.46
U-Net+GradFormer 94.25 97.03 98.83 2.73
U-Net+GFEM 92.32 96.14 96.30 3.96
U-Net+GradFormer+GFEM 94.82 97.34 98.94 1.82
III-C1 Studies of Dilation Rate for GradFormer
As mentioned in II-A, GradFormer employs varying dilation rates across different heads. Tab. I demonstrates the efficacy of this strategy, as models featuring diverse dilation ratios outperform those with a uniform rate. Furthermore, networks employing dilation ratios of 1 and 3 yield superior results compared to those with larger ratios, indicating that smaller dilation ratios are more effective for capturing local differences in infrared small targets, likely due to their small size. Therefore, the dilation ratio significantly impacts target detection, making its configuration crucial for the effectiveness of GradFormer.
TABLE III: Quantitative Comparsion With Different Methods (The best results are bold, second best results are underline.)
Module Metrics NUDT-SIRST IRSTD1K

### v2

# DATransNet: Dynamic Attention Transformer Network for Infrared Small Target Detection

@@ Abstract
Infrared small target detection (ISTD) is widely used in civilian and military applications. However, ISTD encounters several challenges, including the tendency for small and dim targets to be obscured by complex backgrounds. To address this issue, we propose the Dynamic Attention Transformer Network (DATransNet), which aims to extract and preserve edge information of small targets. DATransNet employs the Dynamic Attention Transformer (DATrans), simulating central difference convolutions (CDC) to extract and integrate gradient features with deeper features. Furthermore, we propose a global feature extraction module (GFEM) that offers a comprehensive perspective to prevent the network from focusing solely on details while neglecting the background information. We compare the network with state-of-the-art (SOTA) approaches, and the results demonstrate that our method performs effectively. Our source code is available at https://github.com/greekinRoma/DATransNet.
@@
Infrared small target detection (ISTD), convolution neural network (CNN), Dynamic Attention Transformer, global feature extraction.
@@ I Introduction
INFRARED small target detection (ISTD) is critically important for numerous space-based computer vision applications. Although ISTD has been extensively studied, it remains a challenge. The considerable imaging distances associated with space-based infrared cameras often result in small targets that lack detailed texture and shape information.
@@
Currently, the ISTD methods can be categorized into model-driven and data-driven methods. Model-driven methods include three main approaches. 1) Filter-based methods, such as the Tophat [1], enhance target visibility in complex backgrounds while suppressing background noise. 2) Methods based on the human visual system (HVS), such as the local contrast measure (LCM)[2] and multi-patch contrast measure (MPCM)[3], leverage human visual perception to improve local information. 3) Low-rank matrix decomposition and reconstruction methods, such as infrared patch image (IPI)[4], and Rewight infrared patch tensor (RIPT) [5], decompose small targets and backgrounds using low-rank matrices. However, these techniques require manual parameter settings and often struggle to adapt to complex scenarios.
@@
With the advancement of deep learning, data-driven approaches have achieved substantial progress in infrared small target detection (ISTD). For example, the Asymmetric Context Modulation (ACM) network [6] introduces an asymmetric feature fusion technique, an alternative to conventional skip connections in U-Net. Building on this, DNANet [7] implements a multilayer nested architecture that supports progressive and adaptive interactions among feature layers. Moreover, UIUNet [8] enhances the detection of local target contrasts by integrating multiple U-Net structures and using interactive cross-attention mechanisms for feature fusion. Additionally, GSTUnet [9] merges vision Transformer technology with CNNs in the encoder to efficiently extract edge features from small targets. Generally, these data-driven methods surpass traditional model-driven approaches.
Some challenges faced by data-driven methods, such as the substantial number of parameters and the potential loss of local feature information. Our motivation focus on enhancing the feature information from small targets.
Although Data-driven methods achieve good performance, they suffer from a substantial number of parameters and potential loss of local information. Therefore, we propose some methods to improve the ISTD network, and the new network is called Dynamic Attention Transformer Network (DATransNet). DATransNet incorporates two key innovation components to improve feature extraction. First, we propose the Dynamic Attention Transformer (DATrans), which utilizes Transformers to simulate the central difference convolution (CDC) and enable the network to retain the detailed information of the inputs in deep layers and give dynamic weights to differences result from differenct directions. Second, the global feature extraction module (GFEM) solves the problem of lacking global background perception, thereby improving the ability to obtain contextual information.
The main contributions of this letter are as follows.
@@
    We propose the DATransNet, which utilizes the DATrans to simulate the CDC with dynamic weights, extracting the edge and point information.
@@
    We conduct comparative experiments on the IRSTD-1K and NUDT-SIRST datasets, and the results show that DATransNet outperforms existing methods.
@@ II Methods
The overall architecture of DATransNet is illustrated in Fig. 1. DATrans is based on U-Net and implements DATrans and GFEM to extract detailed information and contextual features.
Figure 1: The overall structure of DATransNet. The red represents the downsampling module and the blue corresponds to the upsampling module.
II-A Dynamic Attention Transformer (DATrans)
CDCs are defined in Eq. 1 and are employed to extract edge and gradient features. Here, $c_{inp}$ represents the number of input channels, $o_{ixy}$ denotes the input value at the $i$-th channel and position $(x,y)$, and $b_{ijxy}$ refers to the $j$-th surrounding value around $o_{ixy}$. $out_{\hat{c}xy}$ denotes the output of the $\hat{c}$-th channel at position $(x,y)$. $w_{\hat{c}ijxy}$ is the weight applied to the difference $(o_{ixy}-b_{ijxy})$, which contributes to determining $out_{\hat{c}xy}$. A significant limitation of CDCs is the use of fixed weights for different directions, meaning the importance of each direction remains constant across different images.
$out_{\hat{c}xy}=\sum^{c_{inp}-1}_{i=0}\sum^{7}_{j=0}w_{\hat{c}ijxy}(b_{ijxy}-o_{ixy})\$ (1)
We can transform Eq. 1 into a fusion of Eq. 2 and Eq. 3. The process described in Eq. 2 is illustrated in Fig. 2. As shown in Fig. 2, we use edge convolution to extract the difference values from the input $T\in\mathbb{R}^{c_{inp}\times w\times h}$ and reshape them to produce a local difference matrix $D\in\mathbb{R}^{8c_{inp}\times wh}$. In Eq. 3, we utilize the weight matrix $M_{w}\in\mathbb{R}^{c_{o}\times 8c_{inp}}$ and the local difference matrix $D$ to obtain the final result $Out$. Here, $w$ denotes the width of the input, $h$ represents the input height, and $c_{o}$ signifies the number of output channels.
@@ II-A Dynamic Attention Transformer (DATrans)
Figure 2: $D$ is obtained by performing edge detection convolution on $T$ and unfolding the output.
When $M_{w}$ is influenced by the input $T$, the CDC could exhibit enhanced robustness. Based on the analysis above, we introduce the DATrans to simulate CDC and extract the details using dynamic weights from low-level layers for ISTD. The structure of DATrans is shown in Fig. 3.
Figure 3: The various heads in DATrans employ the edge convolution to get $D_{\sum}$. Each head’s key and value are derived from $D_{\sum}$, while the query is generated from the flattened input.
In varied heads of the DATrans, we employ different unfolding methods based on edge detection convolutions in Fig. 2 with varying dilation and reshape operations to get the unfolding .
In the various heads of DATrans, we apply different edge detection convolutions from Fig. 2 to $T$, utilizing varying dilation ratios. The outputs from these convolutions are then reshaped into $D^{n}\in\mathbb{R}^{8c_{inp}\times wh}$. Additionally, the input $T$ is reshaped to form $O\in\mathbb{R}^{c_{inp}\times wh}$. The token $O$ serves as the query, and the local difference matrix $D^{n}\in\mathbb{R}^{8c_{inp}\times wh}$ functions as both the key and the value.
@@
where $W^{n}_{Q}\in\mathbb{R}^{\frac{c_{o}}{m}\times c_{inp}}$, $W^{n}_{K}\in\mathbb{R}^{\frac{8c_{inp}}{m}\times 8c_{inp}}$, $W^{n}_{V}\in\mathbb{R}^{\frac{8c_{inp}}{m}\times 8c_{inp}}$ are the reshaped weights of convolutions, $n$ denotes dilation ratios, and $m$ denotes the number of heads. We utilize the $Q^{n}\in\mathbb{R}^{\frac{c_{o}}{m}\times wh}$ and $K^{n}\in\mathbb{R}^{\frac{8c_{inp}}{m}\times wh}$ to produce the attention matrix $M^{n}\in\mathbb{R}^{\frac{c_{o}}{m}\times\frac{8c_{inp}}{m}}$. $V^{n}\in\mathbb{R}^{\frac{8c_{inp}}{m}\times wh}$ denotes the value of the transformer. The result $Out^{n}\in\mathbb{R}^{\frac{c_{o}}{m}\times wh}$ is defined as follows:
@@
where $M^{n}_{mix}\in\mathbb{R}^{\frac{c_{o}}{m}\times 8c_{inp}}$ denotes the dynamic matrix derived from a attention matrix $M^{n}$ and a fixed convolution weight matrix $W_{V}^{n}$. According to Eq. 3, $Out^{n}$ is the equivalent result of the CDC with dynamic weights and a dilation ratio of n. $Norm$ is the Instance Normalization. At last, we concatenate the results in varied heads, reshape them and use $1\times 1$ convolution to fuse them as shown in the Equation below:
@@
II-B Global Feature Extraction Module (GFEM)
Background information is crucial alongside detailed features in ISTD. However, perceiving background relies on global features, which the DATrans struggles to handle effectively. To address this limitation and enhance global information perception, we propose the GFEM, as illustrated in Fig. 4.
@@ II-B Global Feature Extraction Module (GFEM)
Figure 4: The structure of the GFEM using the Non-local (Non-local Attention Module) to capture global spatial features and SE Block (Squeeze-and-Excitation Block) to extract global channel information.
@@ II-C Loss Function
The loss function utilized in our model training is the soft intersection over union (SoftIoU) loss, where the predictions of DATransNet are mapped into the range of 0 and 1, using the Sigmoid function as shown in Eq. 7.
@@
$loss=1-\frac{\sum_{i,j}p_{i,j}\cdot g_{i,j}}{\sum_{i,j}p_{i,j}+\sum_{i,j}g_{i,j}-\sum_{i,j}p_{i,j}\cdot g_{i,j}}$ (8)
@@
Figure 5: Visual results from varied data-driven methods. Red, green, and yellow boxes label detected targets, missed targets, and false alarms. The Red bounding box are zoom in, facilitating to auxiliary our to evaluation detection results. Our DATransNet can achieve high probability detection rate, and achieve precision localization and segmentation of their shapes.
@@ III Experiment and Analysis
This section first describes the experimental details and the evaluation metrics. Following with a series of ablation studies to verify the proposed modules. Finally, a comprehensive comparative experiments with other methods on qualitative and quantitative results demonstrate that our approach outperforms existing state-of-the-art (SOTA) methods.
@@ III-A Dataset and Evaluation Metrics
The datasets used for evaluating the DATransNet are NUDT-SIRST[7] and IRSTD-1K[12]. The evaluation metrics include mean intersection over union ($mIoU$), F-measure ($F_{1}$), the probability of detection ($P_{d}$), and false alarm rate ($F_{a}$). Meanwhile, we plot the ROC curves of various methods to evaluate the robust of DATransNet.
@@ III-B Experimental Environment and Parameter Settings
Our model is implemented by PyTorch framework, then training on NVIDIA GeForce RTX 4080 GPUs. We employ the Adam optimizer, starting with an initial learning rate of $5\times 10^{-4}$. This learning rate is reduced to $5\times 10^{-5}$ at epoch 200 and further decrease to $5\times 10^{-6}$ We use a batch size of 4 for training and run the model for 400 epochs.
@@ III-C Ablation Study
III-C1 Module-wise Performance Gain
In this ablation study, we begin with the most primitive Baseline, Which includes only feature extraction and upsampling with in the U-Net framework. Then, We test the performance of the DATransNet w/o each module, the results are shown in Tab. I. The results illustrate the effectiveness of each module. The performance of our network improves progressively with the inclusion of DATrans and GFEM. DATrans, which focuses on local difference features, enhances the model’s performance more effectively than GFEM, which primarily relies on global perception. In ISTD, the capability to learn gradient information is more critical than access to global context.
TABLE I: Studies Of Different Components on NUDT-SIRST (The best results are bold)
Module mIoU(%) $F_{1}(\%)$ $P_{d}(\%)$ $F_{a}(10^{-6})$
U-Net 91.31 95.44 97.98 4.46
U-Net+DATrans 94.25 97.03 98.83 2.73
U-Net+GFEM 92.32 96.14 96.30 3.96
U-Net+DATrans+GFEM 94.93 97.39 99.04 2.00
III-C2 Studies of Dilation Rate for DATrans
As mentioned in Section II-A, DATrans employs varying dilation rates across different detection heads. Tab. II demonstrates the efficacy of this strategy, as models with diverse dilation ratios outperform those with a uniform rate. Furthermore, networks with dilation ratios of 1 and 3 yield superior results compared to those with other dilation ratios. Therefore, the dilation ratio significantly impacts target detection, and we choose the ratios of 1 and 3 as the dilation ratios for our network.
TABLE II: Studies on different dilation rates on NUDT-SIRST (The best results are bold)
@@ III-C2 Studies of Dilation Rate for DATrans
1,2,3,4 94.58 97.20 98.04 2.21
TABLE III: Studies on GFEM on NUDT-SIRST (The best results are bold)
Non-local SE Block mIoU(%) $F_{1}(\%)$ $P_{d}(\%)$ $F_{a}(10^{-6})$
$\checkmark$ -- 94.69 97.27 98.73 2.57
-- $\checkmark$ 94.53 97.19 98.23 1.75
$\checkmark$ $\checkmark$ 94.93 97.39 99.04 2.00
III-C3 Studies of GFEM
As mentioned in III, we conduct the experiments on the modules in the GFEM. As we can see, the network combining the two modules performs better.
III-D Comparsion with State-of-the-art (SOTA) Methods
To evaluate the effectiveness of our proposed model, we compare it with several SOTA methods on NUDT-SIRST and IRSTD-1K datasets, including SCTransNet [16], UIUNet [8], ACM [6], ALCNet [14], AGPCNet [15], RDIAN [13], and DNANet [7]. As shown in Tab. IV, and our method performs excellently. We utilize DATransNet with dilation ratios of 1 and 5 as the baseline models for comparative evaluation. The comparison experiments include qualitative results and quantitative results, receptively.
III-D1 Qualitative Result
As shown in Fig. 5, our model preserves the shape of targets more closely to ground truth with low missed detections (green boxes) and false alarms (yellow boxes) compared to other methods, achieve satisfied results.
TABLE IV: Quantitative Comparsion With Different Methods on NUDT-SIRST and IRSTD-1K (The best results are bold, second best results are underline.)
Module Metrics NUDT-SIRST IRSTD-1K